Insert seven arithmetic means between 3 and 23.
Answer by AnlytcPhil(1806)

3, __, __, __, __, __, __, __, 23

There are to be 7 terms between 3 and 23.

That means that the whole sequence is to have 9 terms, 
when we count the first term 3 and the last term 23.  
So n = 9.

an = a1 + (n - 1)d

Substitute:

a1 = first term = 3 
n = 9,
an = a9 = last term = 23,

23 = 3 + (9 - 1)d
23 = 3 + 8d

Subtract 3 from both sides:

20 = 8d

Divide both sides by 8

2.5 = d

So the common difference is d = 2.5,
Add 2.5 to the 1st term, getting 3+2.5 = 5.5 for the 2nd term,
Add 2.5 to the 2nd term, getting 5.5+2.5 = 8 for the 3rd term.
etc., etc.,

3, 5.5, 8, 10.5, 13, 15.5, 18, 20.5, 23

Edwin